OpenStudy (anonymous):

Let f(x) = 27x^5 – 33x^4 – 21x^3 and g(x) = 3x2, find the f(x)/g(x)

OpenStudy (anonymous):

HINT: \(\large \frac{a+b+c}{d}=\frac{a}{d}+\frac{b}{d}+\frac{c}{d} \)

OpenStudy (anonymous):

\[ f(x) = 27x^5 – 33x^4 – 21x^3\] \[g(x) = 3x^2\] \[\frac{f(x)}{g(x)}=\frac{27x^5 – 33x^4 – 21x^3}{3x^2}\]

OpenStudy (anonymous):

9x^2 – 11x – 7 9x^2 + 11x + 7 9x^3 – 11x^2 – 7x 9x^3 + 11x^2 + 7x the answer is one of these

OpenStudy (anonymous):

I just gave you the answer

OpenStudy (anonymous):

I suggest factoring out the numerator so you can simplify the solution

OpenStudy (anonymous):

taking x^3 common from numerator \[3x ^{3}*(9x^{2}-11x-7 ) / 3x^{2}\] by cancellation we'll be left with: \[x* (9x^{2}-11x-7)\] multiply x and we'll get \[9x^{3}-11x^{2}-7x\]

OpenStudy (anonymous):

\[\huge \color{green}{\mathbf{Welcome \quad To \quad Openstudy..}}\]

OpenStudy (anonymous):

wow ^

OpenStudy (anonymous):

\[\huge \color{blue} {\checkmark}\]

OpenStudy (anonymous):

^teach me how to do that please

OpenStudy (anonymous):

oh and btw @needthathelp do you understand hoe to solve that question ?!!?

OpenStudy (anonymous):

Go to Latex practicing section you will learn there how to do these things..

OpenStudy (anonymous):

link please ???

OpenStudy (anonymous):

Click on Find More Subjects and select Latex Practicing from there.. See above near the messages and notifications icons it is written there..

OpenStudy (anonymous):

^ now thats time saving :) thanks

OpenStudy (anonymous):

Welcome..